# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v2

@@ E.2 Experiment and Implementation Details
We adopted the AgentMove [^4] https://github.com/tsinghua-fib-lab/agentmove/ library [7], which provides implementations of three LLM methods: LLM-Mob [42], LLM-ZS [1], and LLM-Move [10]. The preprocessing steps used by AgentMove are as follows.
@@
Second, the LLMs are set to return outputs in JSON format, generating the top 5 predicted POI IDs along with an explanation of their reasoning. Following the AgentMove setup and to ensure replicability, we set the generation parameters as follows: a temperature of 0.0, a maximum output length of 1000 tokens, and an input context window capped at 2000 tokens.
@@
All open-source models are quantized using AWQ [25] and served via vLLM [21]. Inference of open-source models was conducted on NVIDIA A100 GPUs provided by SaturnCloud [^8] https://saturncloud.io/ through the NVIDIA Academic Grant Program. We accessed Gemini via the official API and through the Google Cloud Research Credits program. All modified code implementations are publicly available in our main dataset repository.
E.3 Supplementary Results
@@
We provide the full results of our zero-shot POI recommendation results in Table 11 and 12, providing three metrics: Acc@1, Acc@5, and NDCG@5.
@@
Table 11: Performance of zero-shot POI recommendation baselines across 6 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, and Moscow. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
@@
Method Model Beijing Istanbul Jakarta Kuwait City Melbourne Moscow
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.115 0.308 0.226 0.080 0.225 0.160 0.100 0.245 0.174 0.095 0.270 0.185 0.060 0.150 0.108 0.130 0.245 0.187
Qwen 2.5 7B 0.058 0.385 0.218 0.035 0.240 0.148 0.105 0.245 0.179 0.080 0.220 0.155 0.030 0.130 0.083 0.090 0.270 0.185
Llama 3.1 8B 0.000 0.000 0.000 0.020 0.110 0.065 0.055 0.150 0.104 0.030 0.100 0.066 0.010 0.065 0.040 0.030 0.100 0.068
Gemma 2 9B 0.115 0.288 0.206 0.075 0.200 0.146 0.105 0.240 0.178 0.080 0.210 0.150 0.055 0.150 0.108 0.100 0.240 0.176
LLM-ZS Gemini 2 Flash 0.058 0.385 0.246 0.090 0.235 0.166 0.110 0.250 0.188 0.080 0.245 0.167 0.065 0.160 0.115 0.125 0.300 0.217
Qwen 2.5 7B 0.038 0.404 0.237 0.040 0.235 0.141 0.065 0.250 0.161 0.050 0.220 0.140 0.040 0.155 0.100 0.080 0.260 0.176
Llama 3.1 8B 0.077 0.346 0.221 0.040 0.225 0.137 0.045 0.200 0.126 0.060 0.210 0.137 0.040 0.155 0.101 0.080 0.270 0.183
Gemma 2 9B 0.096 0.308 0.217 0.045 0.225 0.141 0.105 0.250 0.180 0.070 0.230 0.153 0.050 0.140 0.100 0.080 0.290 0.194
LLM-Move Gemini 2 Flash 0.096 0.346 0.218 0.205 0.385 0.289 0.295 0.405 0.350 0.220 0.380 0.295 0.225 0.325 0.275 0.220 0.400 0.316
Qwen 2.5 7B 0.192 0.346 0.280 0.175 0.270 0.226 0.115 0.225 0.169 0.160 0.285 0.227 0.110 0.220 0.165 0.230 0.310 0.274
Llama 3.1 8B 0.058 0.135 0.100 0.015 0.055 0.036 0.015 0.025 0.021 0.010 0.035 0.023 0.040 0.195 0.123 0.005 0.065 0.031
Gemma 2 9B 0.096 0.365 0.229 0.100 0.200 0.155 0.235 0.290 0.266 0.120 0.275 0.202 0.115 0.275 0.199 0.110 0.245 0.185
@@
Table 12: Performance of zero-shot POI recommendation baselines across 6 cities: New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
@@
Method Model New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.095 0.175 0.136 0.090 0.220 0.160 0.130 0.305 0.223 0.055 0.160 0.111 0.060 0.160 0.112 0.140 0.320 0.238
Qwen 2.5 7B 0.070 0.185 0.131 0.030 0.195 0.116 0.090 0.290 0.188 0.040 0.170 0.108 0.035 0.145 0.091 0.110 0.350 0.243
Llama 3.1 8B 0.025 0.090 0.061 0.010 0.090 0.050 0.030 0.165 0.098 0.005 0.020 0.013 0.020 0.085 0.053 0.005 0.045 0.025
Gemma 2 9B 0.070 0.175 0.124 0.055 0.185 0.122 0.085 0.230 0.162 0.050 0.150 0.104 0.030 0.130 0.086 0.145 0.345 0.255
LLM-ZS Gemini 2 Flash 0.080 0.170 0.129 0.110 0.210 0.164 0.150 0.315 0.235 0.065 0.160 0.113 0.060 0.155 0.111 0.160 0.380 0.278
Qwen 2.5 7B 0.050 0.180 0.116 0.045 0.175 0.111 0.095 0.290 0.198 0.045 0.155 0.103 0.045 0.170 0.109 0.120 0.365 0.257
Llama 3.1 8B 0.055 0.160 0.111 0.030 0.205 0.123 0.030 0.280 0.159 0.060 0.165 0.116 0.040 0.185 0.110 0.110 0.415 0.269
Gemma 2 9B 0.075 0.185 0.129 0.065 0.185 0.126 0.075 0.300 0.192 0.050 0.165 0.112 0.045 0.155 0.103 0.110 0.395 0.263
LLM-Move Gemini 2 Flash 0.235 0.415 0.325 0.210 0.335 0.273 0.285 0.415 0.350 0.170 0.270 0.221 0.230 0.420 0.331 0.250 0.470 0.368
Qwen 2.5 7B 0.120 0.255 0.188 0.135 0.175 0.155 0.155 0.235 0.199 0.095 0.165 0.133 0.125 0.280 0.205 0.250 0.360 0.312
Llama 3.1 8B 0.035 0.130 0.084 0.040 0.060 0.049 0.045 0.045 0.045 0.020 0.040 0.030 0.055 0.220 0.141 0.030 0.060 0.046
Gemma 2 9B 0.115 0.245 0.183 0.175 0.235 0.208 0.195 0.300 0.252 0.105 0.150 0.128 0.125 0.370 0.254 0.130 0.305 0.225
Appendix F License and Data Usage

### v3

@@ E.3. Experiment and Implementation Details
We adopted the AgentMove [^11] https://github.com/tsinghua-fib-lab/agentmove/ library (Feng et al., 2025), which provides implementations of three LLM methods: LLM-Mob (Wang et al., 2023c), LLM-ZS (Beneduce et al., 2024), and LLM-Move (Feng et al., 2024a). The preprocessing steps used by AgentMove are as follows.
@@
Second, the LLMs are set to return outputs in JSON format, generating the top 5 predicted POI IDs along with an explanation of their reasoning. Following the AgentMove setup and to ensure replicability, we set the generation parameters as follows: a temperature of 0.0, a maximum output length of 1000 tokens, and an input context window capped at 2000 tokens. Following its default configuration, LLM-Move’s candidate set is set to the 100 nearest POIs.
@@
All open-source models are quantized using AWQ (Lin et al., 2024) and served via vLLM (Kwon et al., 2023). Inference of open-source models was conducted on NVIDIA A100 GPUs. We accessed Gemini via the official API. All modified code implementations are publicly available in our main dataset repository.
E.4. Supplementary Results
@@
We provide the full results of our zero-shot POI recommendation results in Table 15, 16, and 17, providing three metrics: Acc@1, Acc@5, and NDCG@5. Additionally, Table 18, 19, and 20 present zero-shot performance across two time periods (2012-2013 and 2017-2018) using LLM-Move.
@@
Table 15. Performance of zero-shot POI recommendation baselines across 5 cities: Bandung, Beijing, Istanbul, Jakarta, Kuwait City. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
@@
Method Model Bandung Beijing Istanbul Jakarta Kuwait City
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.105 0.170 0.139 0.115 0.308 0.226 0.080 0.225 0.160 0.100 0.245 0.174 0.095 0.270 0.185
Qwen 2.5 7B 0.060 0.155 0.111 0.058 0.385 0.218 0.035 0.240 0.148 0.105 0.245 0.179 0.080 0.220 0.155
Llama 3.1 8B 0.010 0.100 0.055 0.000 0.000 0.000 0.020 0.110 0.065 0.055 0.150 0.104 0.030 0.100 0.066
Gemma 2 9B 0.070 0.175 0.126 0.115 0.288 0.206 0.075 0.200 0.146 0.105 0.240 0.178 0.080 0.210 0.150
LLM-ZS Gemini 2 Flash 0.095 0.195 0.147 0.058 0.385 0.246 0.090 0.235 0.166 0.110 0.250 0.188 0.080 0.245 0.167
Qwen 2.5 7B 0.055 0.185 0.126 0.038 0.404 0.237 0.040 0.235 0.141 0.065 0.250 0.161 0.050 0.220 0.140
Llama 3.1 8B 0.045 0.210 0.131 0.077 0.346 0.221 0.040 0.225 0.137 0.045 0.200 0.126 0.060 0.210 0.137
Gemma 2 9B 0.065 0.185 0.130 0.096 0.308 0.217 0.045 0.225 0.141 0.105 0.250 0.180 0.070 0.230 0.153
LLM-Move Gemini 2 Flash 0.225 0.350 0.289 0.096 0.346 0.218 0.205 0.385 0.289 0.295 0.405 0.350 0.220 0.380 0.295
Qwen 2.5 7B 0.100 0.155 0.128 0.192 0.346 0.280 0.175 0.270 0.226 0.115 0.225 0.169 0.160 0.285 0.227
Llama 3.1 8B 0.030 0.035 0.033 0.058 0.135 0.100 0.015 0.055 0.036 0.015 0.025 0.021 0.010 0.035 0.023
Gemma 2 9B 0.175 0.245 0.213 0.096 0.365 0.229 0.100 0.200 0.155 0.235 0.290 0.266 0.120 0.275 0.202
@@
Table 16. Performance of zero-shot POI recommendation baselines across 5 cities: Melbourne, Moscow, New York, Palembang, Petaling Jaya. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
@@
Method Model Melbourne Moscow New York Palembang Petaling Jaya
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.060 0.150 0.108 0.130 0.245 0.187 0.095 0.175 0.136 0.135 0.275 0.208 0.090 0.220 0.160
Qwen 2.5 7B 0.030 0.130 0.083 0.090 0.270 0.185 0.070 0.185 0.131 0.075 0.205 0.143 0.030 0.195 0.116
Llama 3.1 8B 0.010 0.065 0.040 0.030 0.100 0.068 0.025 0.090 0.061 0.005 0.040 0.025 0.010 0.090 0.050
Gemma 2 9B 0.055 0.150 0.108 0.100 0.240 0.176 0.070 0.175 0.124 0.095 0.240 0.171 0.055 0.185 0.122
LLM-ZS Gemini 2 Flash 0.065 0.160 0.115 0.125 0.300 0.217 0.080 0.170 0.129 0.130 0.260 0.196 0.110 0.210 0.164
Qwen 2.5 7B 0.040 0.155 0.100 0.080 0.260 0.176 0.050 0.180 0.116 0.050 0.215 0.135 0.045 0.175 0.111
Llama 3.1 8B 0.040 0.155 0.101 0.080 0.270 0.183 0.055 0.160 0.111 0.070 0.240 0.154 0.030 0.205 0.123
Gemma 2 9B 0.050 0.140 0.100 0.080 0.290 0.194 0.075 0.185 0.129 0.060 0.235 0.150 0.065 0.185 0.126
LLM-Move Gemini 2 Flash 0.225 0.325 0.275 0.220 0.400 0.316 0.235 0.415 0.325 0.260 0.385 0.329 0.210 0.335 0.273
Qwen 2.5 7B 0.110 0.220 0.165 0.230 0.310 0.274 0.120 0.255 0.188 0.130 0.195 0.163 0.135 0.175 0.155
Llama 3.1 8B 0.040 0.195 0.123 0.005 0.065 0.031 0.035 0.130 0.084 0.010 0.015 0.013 0.040 0.060 0.049
Gemma 2 9B 0.115 0.275 0.199 0.110 0.245 0.185 0.115 0.245 0.183 0.210 0.270 0.240 0.175 0.235 0.208
Table 17. Performance of zero-shot POI recommendation baselines across 5 cities: São Paulo, Shanghai, Sydney, Tangerang, Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Method Model São Paulo Shanghai Sydney Tangerang Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.130 0.305 0.223 0.055 0.160 0.111 0.060 0.160 0.112 0.155 0.285 0.225 0.140 0.320 0.238
Qwen 2.5 7B 0.090 0.290 0.188 0.040 0.170 0.108 0.035 0.145 0.091 0.095 0.285 0.196 0.110 0.350 0.243
Llama 3.1 8B 0.030 0.165 0.098 0.005 0.020 0.013 0.020 0.085 0.053 0.020 0.120 0.073 0.005 0.045 0.025
Gemma 2 9B 0.085 0.230 0.162 0.050 0.150 0.104 0.030 0.130 0.086 0.145 0.270 0.209 0.145 0.345 0.255
LLM-ZS Gemini 2 Flash 0.150 0.315 0.235 0.065 0.160 0.113 0.060 0.155 0.111 0.145 0.310 0.234 0.160 0.380 0.278
Qwen 2.5 7B 0.095 0.290 0.198 0.045 0.155 0.103 0.045 0.170 0.109 0.100 0.315 0.215 0.120 0.365 0.257
Llama 3.1 8B 0.030 0.280 0.159 0.060 0.165 0.116 0.040 0.185 0.110 0.080 0.255 0.173 0.110 0.415 0.269
Gemma 2 9B 0.075 0.300 0.192 0.050 0.165 0.112 0.045 0.155 0.103 0.100 0.330 0.227 0.110 0.395 0.263
LLM-Move Gemini 2 Flash 0.285 0.415 0.350 0.170 0.270 0.221 0.230 0.420 0.331 0.200 0.340 0.274 0.250 0.470 0.368
Qwen 2.5 7B 0.155 0.235 0.199 0.095 0.165 0.133 0.125 0.280 0.205 0.175 0.280 0.229 0.250 0.360 0.312
Llama 3.1 8B 0.045 0.045 0.045 0.020 0.040 0.030 0.055 0.220 0.141 0.000 0.005 0.003 0.030 0.060 0.046
Gemma 2 9B 0.195 0.300 0.252 0.105 0.150 0.128 0.125 0.370 0.254 0.125 0.250 0.193 0.130 0.305 0.225
Table 18. Performance of zero-shot POI recommendation using LLM-Move across two time periods and 5 cities: Bandung, Beijing, Istanbul, Jakarta, and Kuwait City. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Time Period Model Bandung Beijing Istanbul Jakarta Kuwait City
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
2012-2013 Gemini 2 Flash 0.227 0.351 0.290 0.102 0.367 0.232 0.212 0.384 0.290 0.295 0.409 0.352 0.423 0.500 0.453
Qwen 2.5 7B 0.098 0.155 0.126 0.204 0.367 0.298 0.192 0.295 0.247 0.114 0.223 0.167 0.269 0.423 0.357
Llama 3.1 8B 0.031 0.036 0.034 0.041 0.122 0.086 0.007 0.027 0.017 0.010 0.021 0.016 0.000 0.000 0.000
Gemma 2 9B 0.180 0.247 0.217 0.102 0.388 0.244 0.116 0.199 0.159 0.228 0.285 0.260 0.308 0.385 0.342
2017-2018 Gemini 2 Flash 0.167 0.333 0.272 0.000 0.000 0.000 0.185 0.389 0.284 0.286 0.286 0.286 0.190 0.362 0.271
Qwen 2.5 7B 0.167 0.167 0.167 0.000 0.000 0.000 0.130 0.204 0.168 0.143 0.286 0.233 0.144 0.264 0.208
Llama 3.1 8B 0.000 0.000 0.000 0.333 0.333 0.333 0.037 0.130 0.088 0.143 0.143 0.143 0.011 0.040 0.026
Gemma 2 9B 0.000 0.167 0.083 0.000 0.000 0.000 0.056 0.204 0.142 0.429 0.429 0.429 0.092 0.259 0.181
Table 19. Performance of zero-shot POI recommendation using LLM-Move across two time periods and 5 cities: Melbourne, Moscow, New York, Palembang, Petaling Jaya. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Time Period Model Melbourne Moscow New York Palembang Petaling Jaya
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
2012-2013 Gemini 2 Flash 0.226 0.329 0.279 0.218 0.401 0.316 0.240 0.403 0.320 0.256 0.385 0.327 0.199 0.348 0.274
Qwen 2.5 7B 0.116 0.232 0.175 0.234 0.310 0.275 0.130 0.240 0.190 0.128 0.195 0.162 0.142 0.184 0.164
Llama 3.1 8B 0.039 0.200 0.125 0.005 0.066 0.032 0.032 0.117 0.074 0.010 0.015 0.013 0.014 0.043 0.028
Gemma 2 9B 0.097 0.271 0.189 0.112 0.249 0.188 0.130 0.260 0.198 0.215 0.272 0.244 0.199 0.262 0.234
2017-2018 Gemini 2 Flash 0.222 0.311 0.261 0.333 0.333 0.333 0.217 0.457 0.342 0.400 0.400 0.400 0.237 0.305 0.273
Qwen 2.5 7B 0.089 0.178 0.132 0.000 0.333 0.167 0.087 0.304 0.183 0.200 0.200 0.200 0.119 0.153 0.134
Llama 3.1 8B 0.044 0.178 0.116 0.000 0.000 0.000 0.043 0.174 0.116 0.000 0.000 0.000 0.102 0.102 0.102
Gemma 2 9B 0.178 0.289 0.232 0.000 0.000 0.000 0.065 0.196 0.133 0.000 0.200 0.086 0.119 0.169 0.144
Table 20. Performance of zero-shot POI recommendation using LLM-Move across two time periods and 5 cities: São Paulo, Shanghai, Sydney, Tangerang, Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Time Period Model São Paulo Shanghai Sydney Tangerang Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
2012-2013 Gemini 2 Flash 0.298 0.440 0.369 0.192 0.288 0.242 0.256 0.462 0.367 0.197 0.338 0.272 N/A N/A N/A
Qwen 2.5 7B 0.173 0.250 0.215 0.109 0.186 0.151 0.122 0.288 0.209 0.172 0.278 0.226 N/A N/A N/A
Llama 3.1 8B 0.048 0.048 0.048 0.006 0.026 0.016 0.064 0.224 0.148 0.000 0.005 0.003 N/A N/A N/A
Gemma 2 9B 0.202 0.315 0.264 0.109 0.160 0.136 0.122 0.378 0.257 0.126 0.253 0.195 N/A N/A N/A
2017-2018 Gemini 2 Flash 0.219 0.281 0.251 0.091 0.205 0.147 0.136 0.273 0.204 0.500 0.500 0.500 0.250 0.470 0.368
Qwen 2.5 7B 0.063 0.156 0.115 0.045 0.091 0.070 0.136 0.250 0.192 0.500 0.500 0.500 0.250 0.360 0.312
Llama 3.1 8B 0.031 0.031 0.031 0.068 0.091 0.077 0.023 0.205 0.117 0.000 0.000 0.000 0.030 0.060 0.046
Gemma 2 9B 0.156 0.219 0.189 0.091 0.114 0.101 0.136 0.341 0.246 0.000 0.000 0.000 0.130 0.305 0.225
Appendix F Spatiotemporal Classification and Reasoning: Task Details
F.1. Problem Formulation
Borrowing the notation used in Section C, we formulate this task as follows. Given a current contextual trajectory $T^{\prime}_{u}(t)$ of user $u$ up to time $t$, the goal of spatiotemporal trajectory classification is to predict a property $y$ of the trajectory. In this study, we focus on weekday/weekend classification, where $y\in\{\text{weekday},\text{weekend}\}$.
Formally, the LLM serves as a classification function:
$f:T^{\prime}_{u}(t)\rightarrow\hat{y}$
where $\hat{y}$ denotes the predicted class label for the trajectory. The model is evaluated based on its accuracy in correctly classifying trajectories according to this property.
F.2. Experiment and Implementation Details
Preprocessing
We borrowed the experimental setup of AgentMove, similar to our zero-shot POI recommendation procedure in Section E.3. That is, we selected 200 random users from the test set and sampled one random trajectory for each user. This trajectory is then included in the test set. Each check-in is described by four attributes: the hour (in 12-hour format), the day of the week, the POI ID, and the POI category name.
LLMs are set to return outputs in a structured/JSON format, predicting whether the trajectory ended on a weekday or a weekend, along with an explanation of their reasoning. To ensure replicability, Gemini and GPT-4 models are set with the following generation parameters: a temperature of 0.0, a maximum output length of 1000 tokens, and an input context window capped at 2000 tokens. Due to API requirements, GPT-5 Nano uses the following generation parameters: a fixed temperature of 1.0, a maximum output length of 4096 tokens, low verbosity, and medium reasoning effort.
Prompt
Prompt template for spatiotemporal weekday-weekend classification is shown in Listing 4.
1A trajectory is a sequence of check-ins, each represented as (start_time, poi_category). The detailed explanation of each element is as follows:
2start_time: the start time of the check-in in 12h clock format.
3poi_category: the category of the point of interest (POI) visited during the check-in
4
5The trajectory is as follows: {[check-in time-of-day, POI category] for check-in in trajectory}
6
7Your task is to classify whether the last check-in occurs on a weekday or a weekend.
8Consider the temporal information (i.e., start_time) of the trajectory, which is important because peoples activity varies during different time (e.g., nighttime versus daytime).
9Consider the POI categories, which can provide insights into the users activity patterns.
10Also consider the city context, as different cities may have different cultural and social norms that influence activity patterns. The city is: {city}.
11
12Please organize your answer in a JSON object containing following keys:
13"prediction" ("weekday" or "weekend") and "reason" (a concise explanation that supports your prediction).
14Do not include line breaks in your output.
Listing 4: Prompt for Weekday vs. Weekend Classification
Models and Implementations
We use the following LLMs in our experiments:
  • Gemini 2.0 Flash (gemini-2.0-flash),
  • GPT-4o Mini (gpt-4o-mini),
  • GPT-4.1 Mini (gpt-4.1-mini),
  • GPT-5 Nano (gpt-5-nano).
We accessed Gemini and GPT models via the official API. All modified code implementations are publicly available in our main dataset repository.
Appendix G License and Data Usage